# Investigating the Kinetic Effects on Current Gradient-Driven Instabilities of Electron Current Layers via Particle-in-Cell Simulations

### v4

@@ 1 Introduction
Plasma is a state of matter in which electrons and ions move, exhibiting collective behavior under the influence of electric and magnetic fields generated either by their own motion or by external sources. Magnetic fields significantly influence plasma dynamics, leading to rich and complex behaviors. They affect the motion of charged particles, which, in turn, modify the magnetic fields due to the “frozen-in” condition arising from plasma’s high electrical conductivity. A key phenomenon where the magnetic field is altered by plasma particles is magnetic reconnection (MR). In MR, plasma pushes two oppositely directed magnetic field lines towards each other, changing the topology of the magnetic field lines [1, 2]. MR occurs in various natural settings, including within the solar system [3]. It plays a crucial role in solar flares, coronal mass ejections, the Earth’s magnetosphere, and astrophysical jets, and is also believed to occur during star formation. Additionally, MR is observed in several fusion plasma devices, such as Tokamak, Reverse Field Pinch, and Spheromak configurations [4, 5]. Numerous laboratory plasma devices have been developed to investigate the mechanisms behind MR [6, 7, 8].
Significant efforts have been made to propose theoretical models that explain MR at different time and length scales in various plasma conditions. The resistive MHD (magnetohydrodynamics) models, such as the Sweet-Parker and Petschek models, describe MR occurring in long, narrow current sheets with a reconnection rate limited by plasma resistivity. The Sweet-Parker model [9, 10] predicts a slow reconnection rate, insufficient to explain the fast reconnection events observed in nature. In contrast, the Petschek model [11] suggests a faster reconnection process, although it is less applicable to collisionless plasmas. In space and astrophysics contexts, where fast reconnection occurs, plasmas are known to be collisionless. In a collisionless plasma, where the mean free path of particles is much larger than the system size, reconnection dynamics are predominantly governed by electromagnetic forces rather than collisional interactions.
@@
Gaur and Kaw [13] investigated the susceptibility of thin electron current sheets to tearing and surface-preserving instabilities within 2D EMHD when equilibrium length scales are shorter than the ion skin depth. These instabilities, driven by current gradients, exhibit distinct behaviors: the tearing mode results in magnetic island formation, while the surface-preserving mode maintains magnetic flux surfaces. The tearing mode is non-local, unstable to larger-scale perturbations, while the surface-preserving mode is local, unstable to shorter wavelength perturbations. Similar studies by Lukin [19] and Jain et al. [20, 21] have examined these modes within the EMHD framework.
@@
2 Particle-in-cell (PIC) method and Simulation Setup
Particle simulation model can be classified into three principal types: the particle-particle (PP) model, the particle-mesh (PM) model, and the particle-particle-particle-mesh (PPPM or P3M) model [23]. The PP model utilizes the action-at-a-distance approach of the force law, the PM model implements the force as a field quantity (approximating it on a mesh), and the P3M model incorporates characteristics of both the PP and PM models.
The Particle-in-Cell (PIC) method is based on the PM model. In PIC simulations, the plasma is represented by a large number of superparticles, each representing a cluster of real particles, typically electrons and ions. These superparticles move through a computational grid (mesh) where the electromagnetic fields are calculated. The position and velocity of each particle are given as initial conditions, which determine the continuous distribution of the particles’ position and velocity in space.
The PIC algorithm involves a self-consistent cycle: assigning charge to the mesh, solving Maxwell’s equations and computing the fields from the mesh-defined potential, and updating the particle velocity and position based on the fields. Fig. 1 shows the algorithm for the PIC technique, which includes the Boris scheme for pushing particles and a finite difference solver for the field calculations . Various PIC codes are available for plasma simulations, with OSIRIS being one of the most widely used.
Figure 1: Overview of the PIC Algorithm Workflow
OSIRIS is a state-of-the-art PIC simulation code that is fully relativistic, massively parallel and fully object-oriented. Written in Fortran-90, it is designed for high-performance computing with multi-dimensional capabilities [24]. OSIRIS simulates the behavior of plasmas by solving the Vlasov-Maxwell equations using the PIC method and is optimized for large-scale parallel supercomputers, utilizing the Message Passing Interface (MPI) for communication between processors [25].
Widely used in plasma physics research, OSIRIS features comprehensive physics models, including electromagnetic fields, relativistic particle dynamics, collisions, and radiation mechanisms. It excels in modeling complex plasma interactions, such as laser-plasma interactions for inertial confinement fusion [26] and particle acceleration in astrophysical phenomena like magnetic reconnection, gamma-ray bursts, and relativistic jets [27].
We have utilized the OSIRIS 4.0 framework [28] to model the 2D tearing and surface preserving instabilities. In our simulations, the normalized units are defined as follows: length is normalized by electron skin depth c/$\omega_{pe}$, time is normalized by $\omega_{pe}^{-1}$, and magnetic field by the quantity $\frac{m_{e}c^{2}}{e(c/\omega_{pe})}$. Where c is the speed of light, $\omega_{pe}$ is the electron plasma frequency, and $m_{e}$ is the electron mass.
@@ 2 Particle-in-cell (PIC) method and Simulation Setup
In all our simulation runs, we have chosen $B_{00}=0.1$ and $\epsilon=1$. A 2D rectangular simulation box with dimensions $L_{x}=8(c/\omega_{pe})$ and $L_{y}=3(c/\omega_{pe})$ has been chosen. The spatial resolution corresponds to a grid size of $\Delta x=2L_{x}/512$ and $\Delta y=2L_{y}/256$. A uniform plasma number density $n_{0}=10^{18}cm^{-3}$ has been chosen. We allow the equilibrium to evolve against the numerical perturbations. We have taken a periodic boundary conditions for both fields and particles because the equilibrium profile chosen is such that it ceases at both the boundaries. The equilibrium magnetic field and the corresponding equilibrium velocity profiles are shown in Fig. 2. Three different magnetic field profiles corresponds to three different values of $C_{0}$ viz., 0, 0.025, 0.06.
@@ 3.1 $C_{0}=0$: Only Tearing Mode Case
We begin by setting $T=0$, corresponding to the typical tearing instability case, which has been extensively studied by various authors using both fluid and PIC simulations. We reproduced this scenario to validate our simulation setup. In Fig. 3, we plot the evolution of the perturbation energy over time. During the initial phase of the simulation, the perturbation energy grows exponentially. The slope of this portion (a straight line on the semilog plot) closely matches twice the maximum growth rate $2\gamma$ derived from linear instability calculations [13]. The dashed line alongside the simulation curve has a slope equal to $2\gamma$, showing good agreement. As the amplitude of the perturbed field increases, nonlinear effects become significant, leading to the saturation of the perturbation energy in the later stages of the simulation.
@@
Figure 4: Shaded isocontours of out-of-plane magnetic field with superimposed streamlines (solid lines) over various times for pure tearing mode case ($C_{0}=0$) and T = 0 eV case.
@@
We now add $C_{0}$, a uniform magnetic field along the direction of equilibrium magnetic field. We choose $C_{0}=0.06$, as this would make the condition $B_{0}-B_{0}^{{}^{\prime\prime}}\leq 0$ for the existence of the tearing mode to be violated. Coincidentally, the opposite condition $B_{0}-B_{0}^{{}^{\prime\prime}}>0$ favors the existence of the surface preserving instability. Hence, only the surface preserving mode will be present in this case [13].
Simulation run with $T=0~{}eV$, show no formation of quadrupoles or islands, which is unlike the tearing mode scenario. Magnetic field lines in this case do not bend and instead form a channel-like structure, as depicted in Fig. 7. In order to understand the behaviour of growth rate of surface preserving mode we conduct simulation runs for the following temperature values: $(T=5,10,20,50,100,200,500)$ electron volts. In Fig. 8, we plot the growth rate vs temperature. In this case, the growth rate goes on to increase, unlike the tearing mode case where the growth rate first increases and then start decreasing after a certain temperature. This can be explained as follows. Surface preserving mode, also termed as stationary nontearing inertial scale mode, is driven by the electron flow perturbations along the direction of the current gradient[19]. An increase in temperature increases the electron flow velocity which, in turn, makes the instability more favourable.
Figure 7: Shaded isocontours of out-of-plane magnetic field with superimposed streamlines (solid lines) over various times for the only surface preserving mode case ($C_{0}=0.06$) and T = 0 eV case.
@@
Figure 9: Shaded isocontours of out-of-plane magnetic field with superimposed streamlines (solid lines) over various times for mixed-modes case ($C_{0}=0.025$) and T = 0 eV case.
Figure 10: Shaded isocontours of the current $J_{z}$ with superimposed streamlines (black lines) over various times for mixed-modes case ($C_{0}=0.025$) and T = 0 eV case.
Figure 11: Shaded isocontours of the current $J_{z}$ with superimposed streamlines (black lines) over various times for pure tearing mode case ($C_{0}=0$) and T = 0 eV case.
@@ 4 Summary and Future Scope
To study the surface-preserving instability, we added a uniform magnetic field $C_{0}=0.06$ along the equilibrium magnetic field direction, which inhibits the tearing mode in the system. To analyze the effect of temperature on the growth rate of this mode, we ran simulations at various temperatures. The growth rate versus temperature plot shows a consistent rise, which can be attributed to the increased electron flow velocity at higher temperatures. This rise favors the instability, as it is driven by perturbations in the flow.

### v5

@@ 1 Introduction
Plasma is a state of matter in which electrons and ions move, exhibiting collective behavior under the influence of electric and magnetic fields generated either by their own motion or by external sources. Magnetic fields significantly influence plasma dynamics, leading to rich and complex behaviors. They affect the motion of charged particles, which, in turn, modify the magnetic fields due to the “frozen-in” condition arising from plasma’s high electrical conductivity. A key phenomenon where the magnetic field is altered by plasma particles is magnetic reconnection (MR). In MR, plasma pushes two oppositely directed magnetic field lines towards each other, changing the topology of the magnetic field lines [1, 2]. MR occurs in various natural settings, including within the solar system [3]. It plays a crucial role in solar flares, coronal mass ejections, the Earth’s magnetosphere, and astrophysical jets, and is also believed to occur during star formation. Additionally, MR is observed in several fusion plasma devices, such as Tokamak [4], Reverse Field Pinch [5], and Spheromak configurations [6, 7]. Numerous laboratory plasma devices have been developed to investigate the mechanisms behind MR [8, 9, 10].
Significant efforts have been made to propose theoretical models that explain MR at different time and length scales in various plasma conditions. The resistive MHD (magnetohydrodynamics) models, such as the Sweet-Parker and Petschek models[11], describe MR occurring in long, narrow current sheets with a reconnection rate limited by plasma resistivity. The Sweet-Parker model predicts a slow reconnection rate, insufficient to explain the fast reconnection events observed in nature. In contrast, the Petschek model suggests a faster reconnection process, although it is less applicable to collisionless plasmas[3]. In space and astrophysics contexts, where fast reconnection occurs, plasmas are known to be collisionless. In a collisionless plasma, where the mean free path of particles is much larger than the system size, reconnection dynamics are predominantly governed by electromagnetic forces rather than collisional interactions.
@@
Gaur and Kaw [13] investigated the susceptibility of thin electron current sheets to tearing and surface-preserving instabilities within 2D EMHD when equilibrium length scales are shorter than the ion skin depth. These instabilities, driven by current gradients [^2] For a discussion on the role of current gradient, the reader is referred to the Appendix., exhibit distinct behaviors: the tearing mode results in magnetic island formation, while the surface-preserving mode maintains magnetic flux surfaces. The tearing mode is non-local, unstable to larger-scale perturbations, while the surface-preserving mode is local, unstable to shorter wavelength perturbations. Similar studies by Lukin [19] and Jain et al. [20, 21] have examined these modes within the EMHD framework.
@@
2 Simulation Setup
Various PIC codes [23] are available for plasma simulations, with OSIRIS being one of the most widely used. OSIRIS is a state-of-the-art PIC simulation code that is fully relativistic, massively parallel and fully object-oriented. Written in Fortran-90, it is designed for high-performance computing with multi-dimensional capabilities [24].
We have utilized the OSIRIS 4.0 framework [25, 26, 27, 28] to model the 2D tearing and surface preserving instabilities. In our simulations, the normalized units are defined as follows: length is normalized by electron skin depth c/$\omega_{pe}$, time is normalized by $\omega_{pe}^{-1}$, velocity is normalised by c and magnetic field by the quantity $\frac{m_{e}c^{2}}{e(c/\omega_{pe})}$. Where c is the speed of light, $\omega_{pe}$ is the electron plasma frequency, and $m_{e}$ is the electron mass. In this study, all physical quantities are expressed in cgs units.
In our simulations, we use a Cartesian coordinate system where the axes are defined as follows: the x-axis represents the shear or gradient direction, corresponding to the current sheet gradient; the y-axis aligns with the equilibrium magnetic field (Beq), representing a uniform in-plane magnetic field; and the z-axis corresponds to the symmetry or out-of-plane direction. This z-axis is associated with equilibrium velocity/current density, and also serves as the direction of the applied guide magnetic field $B_{G}$. The coordinate system and field orientations are depicted schematically in Fig. 1.
Figure 1: Schematic representation of the coordinate system and field orientations in the simulation setup.
@@ 2 Simulation Setup
In all our simulation runs, we have chosen $B_{00}=0.1$ and $\epsilon=1$. A 2D rectangular simulation box with dimensions $L_{x}=8(c/\omega_{pe})$ and $L_{y}=3(c/\omega_{pe})$ has been chosen. The spatial resolution corresponds to a grid size of $\Delta x=2L_{x}/512$ and $\Delta y=2L_{y}/256$. The total number of particles per cell is 50. We allow the equilibrium to evolve against the numerical perturbations. We have taken a periodic boundary conditions for both fields and particles because the equilibrium profile chosen is such that it ceases at both the boundaries. The equilibrium magnetic field and the corresponding equilibrium velocity profiles are shown in Fig. 2. Three different magnetic field profiles correspond to three different values of $C_{0}$ viz., 0, 0.025, 0.06.
@@ 3.1 $C_{0}=0$: Only Tearing Mode Case
We begin by setting $T=0$, corresponding to the typical tearing instability case, which has been extensively studied by various authors using both fluid and PIC simulations. We reproduced this scenario to validate our simulation setup. In Fig. 3, we plot the evolution of the perturbation energy over time. During the initial phase of the simulation, the perturbation energy grows exponentially. The slope of this portion (a straight line on the semilog plot) closely matches twice the maximum growth rate $2\gamma$ derived from linear instability calculations [13]. The dashed line alongside the simulation curve has a slope equal to $2\gamma$, showing good agreement. As the amplitude of the perturbed field increases, nonlinear effects become significant, leading to saturation of the perturbation energy in the later stages of the simulation.
@@
Figure 4: Shaded isocontours of out-of-plane magnetic field ($B_{z}$) with superimposed streamlines (solid lines) over various times for pure tearing mode case ($C_{0}=0$) and T = 0 eV case. The color bar indicates the normalized amplitude of $B_{z}$.
@@
We now add $C_{0}$, a uniform magnetic field aligned with the equilibrium magnetic field. Here, $B_{0}$ represents the equilibrium magnetic field, and $B_{0}^{\prime\prime}=d^{2}B_{0}/dx^{2}$ is its second derivative.The quantity $B_{0}-B_{0}^{\prime\prime}$ must change sign for instabilities to exist, a necessary condition highlighting the role of the current gradient. The condition $B_{0}B_{0}^{\prime\prime}\leq 0$ is a well-known criterion for the tearing instability, as discussed in the literature [12]. Conversely, the condition $B_{0}B_{0}^{\prime\prime}>0$ corresponds to the existence of the surface-preserving mode, as detailed by Gaur and Kaw [13].
In our setup, we choose $C_{0}=0.06$, which ensures that the condition $B_{0}B_{0}^{\prime\prime}\leq 0$ for tearing instability is violated. At the same time, the condition $B_{0}B_{0}^{\prime\prime}>0$ is satisfied, favoring the surface-preserving mode. As a result, only the surface-preserving mode is present in this configuration.
Simulation run with $T=0~{}eV$, show no formation of quadrupoles or islands, which is unlike the tearing mode scenario. Magnetic field lines in this case do not bend and instead form a channel-like structure, as depicted in Fig. 7. In order to understand the behaviour of growth rate of surface preserving mode we conduct simulation runs for the following temperature values: $(T=5,10,20,50,100,200,500)$ electron volts. In Fig. 8, we plot the growth rate vs temperature. In this case, the growth rate goes on to increase, unlike the tearing mode case where the growth rate first increases and then start decreasing after a certain temperature. This can be explained as follows. Surface preserving mode, also termed as stationary nontearing inertial scale mode, is driven by the electron flow perturbations along the direction of the current gradient[19]. An increase in temperature increases the thermal velocity of electrons, which enhances their mobility and indirectly amplifies the electron flow perturbations, making the instability more favorable.
Figure 7: Shaded isocontours of out-of-plane magnetic $B_{z}$ field with superimposed streamlines (solid lines) over various times for the only surface preserving mode case ($C_{0}=0.06$) and T = 0 eV case. The color bar indicates the normalized amplitude of $B_{z}$.
@@
Figure 9: Shaded isocontours of out-of-plane magnetic field $B_{z}$ with superimposed streamlines (solid lines) over various times for mixed-modes case ($C_{0}=0.025$) and T = 0 eV case. The color bar indicates the normalized amplitude of $B_{z}$.
Figure 10: Shaded isocontours of the current $J_{z}$ with superimposed streamlines (black lines) over various times for mixed-modes case ($C_{0}=0.025$) and T = 0 eV case. The color bar indicates the normalized amplitude of $J_{z}$.
Figure 11: Shaded isocontours of the current $J_{z}$ with superimposed streamlines (black lines) over various times for pure tearing mode case ($C_{0}=0$) and T = 0 eV case. The color bar indicates the normalized amplitude of $J_{z}$.
@@ 4 Summary and Future Scope
To study the surface-preserving instability, we added a uniform magnetic field $C_{0}=0.06$ along the equilibrium magnetic field direction, which inhibits the tearing mode in the system. To analyze the effect of temperature on the growth rate of this mode, we ran simulations at various temperatures. The growth rate versus temperature plot shows a consistent rise, which can be attributed to the increased electron flow perturbations at higher temperatures. This rise favors the instability, as it is driven by perturbations in the flow.
@@
  • To focus on electron dynamics, we approximated the ions as stationary, thereby neglecting ion motion. While this approach simplifies the analysis, future work will incorporate mobile ions to explore their dynamic effects, such as the electron-cyclotron drift instability and the interplay with additional magnetic field components.
@@
4.1 Acknowledgments
SM gratefully acknowledges the financial support provided by the Govt. of Gujarat under SHODH scheme. SM and GG would like to thank Professor Abhijit Sen for useful discussions. GG would like to thank Jugal Chowdhury for providing the comments on the manuscript.
Appendix: Role of Current Gradient
The current gradient serves as the free energy source driving tearing and surface-preserving instabilities.
In the absence of ion motion (i.e., $v_{i}=0$), the current density is:
$\mathbf{J}=-n_{e}e\mathbf{v}_{e},$
If $n_{e}$ is uniform, as in our case, the gradient in the equilibrium current becomes:
$J_{e}^{\prime}=\frac{dJ_{e}}{dx}=-n_{e}ev_{e}^{\prime},$
where $v_{e}^{\prime}$ is the spatial gradient in the equilibrium electron velocity.
For a one-dimensional current configuration, the current gradient $J_{e}^{\prime}(x)$ is inversely proportional to the spatial scale of current variation, as expressed by:
$J_{e}^{\prime}=\frac{\Delta J_{0}}{\epsilon},$
where $\epsilon$ represents the spatial scale. Thin current configurations (such as current sheets) represent sharper current gradients, which are critical in driving instabilities. Maxwell’s relation gives:
$J_{e}^{\prime}=B_{0}^{\prime\prime}\left(=\frac{d^{2}B_{0}}{dx^{2}}\right)=-n_{e}ev_{e}^{\prime},$
The one-dimensional equilibrium magnetic field, corresponding to a thin current sheet, aligns in opposite directions across the current sheet. Non-ideal effects such as resistivity, electron inertia, and turbulence cause the magnetic field to decay, inducing plasma motion to fill the space created by the decaying fields.
The quantity $B_{0}-B_{0}^{\prime\prime}$ (or equivalently $B_{0}+v_{e}^{\prime}$) must change sign for instabilities to exist, a necessary condition highlighting the role of the current gradient [13].